OpenStudy (anonymous):

Let f(x)=-5x-1, h(x)=5/x-5 Find (f * h)(0)

OpenStudy (anonymous):

h(0)=5/(0-5) h(0)=5/(-5) h(0)=-1 f(-1)=-5*-1-1 f(-1)=4 therefore (f*h)(0)=4

OpenStudy (anonymous):

is this \[f\circ h(0)\]

OpenStudy (anonymous):

if so shadow has it

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

did you like \[f\circ h(0)\]?

OpenStudy (anonymous):

the answer was 4

OpenStudy (anonymous):

why won't it let me give shadow a medal?

OpenStudy (anonymous):

satellite, how did you make that circle operator?